# KA-TP-05-2025

### v1

@@ 5.4 Boltzmann Equation
After the computation of the TAC, we can move on to the solution of Eq. (5.1). For the effective degrees of freedom $g_{\text{eff}}$, $h_{\text{eff}}$ and $g_{*}^{1/2}$ we use interpolated values from [54]. We determine the freeze-out point $x_{\text{f}}$ via the condition given in [54],
@@
where $\delta$ can be freely chosen. To find the point $x_{\text{f}}$ which fulfills Eq. (5.13), we use a Bisection method [^9] We only search between $x=5$ and $x=50$. If a point is not found within these boundaries it is not a valid freeze-out point and we return 0 for the relic density. to find the root of the equation. If the option Fast is set to true we use $\delta=1.5$ as recommended in [54] and compute the yield via the freeze-out approximation, i.e.
@@
where the $Y_{\text{eq}}^{2}$ term in Eq. (5.1) was discarded. As stated in [17] and [54], this approximation is accurate within a few percent in almost all scenarios, which we can confirm. When the Fast option is turned off we use $\delta=0.1$ to find the freeze-out point. With this we solve Eq. (5.1) numerically with the initial condition $Y_{\text{f}}=1.1Y_{\text{eq}}(x_{\text{f}})$ at the freeze-out point using an adaptive Dormand-Prince 853 method as described in [60].
@@ 6 Validation
In Fig. 2, we present the relic density $\Omega_{c}h^{2}$ as a function of the DM mass $m_{\text{DM}}$ for the real singlet model [62]. The Higgs sector of this model features a DM particle and a SM-like Higgs with mass 125 GeV. For the plot, we chose the mass range of the DM particle between 5 GeV and 1000 GeV, and the freeze-out portal coupling was set to 0.1 (orange and green points) and 6 (red and black points). We observe, as expected, a resonant annihilation at $m_{\text{DM}}=m_{h}/2$, where $m_{h}$ denotes the SM-like Higgs mass value of 125 GeV, resulting in a dip in the relic density. The subsequent peaks and dips in the relic density are a consequence of the DM mass passing the thresholds for on-shell production of, respectively, a $W$ boson, $Z$ boson and finally a SM-like Higgs pair. The dip at the Higgs-mass threshold is much less pronounced for $\lambda_{\text{portal}}=0.1$ than for $\lambda_{\text{portal}}=6$, reflecting the different values of the SM-like Higgs coupling to the DM particles in the two chosen scenarios. As can be inferred from the figure, the MicrOMEGAs and the RelExt results are in very good agreement, with differences of at most 6%.
@@
Next we compare results in the model CP in the Dark [43, 44]. It is based on an N2HDM with an imposed discrete symmetry such that the Higgs spectrum consists of a SM Higgs boson and the additional Higgs bosons, three neutral and two charged ones, residing in a dark sector with explicit CP violation. The lightest of these dark scalars must be neutral and represents the DM candidate. In Fig. 3, we show the absolute value of the relative difference in the relic density between RelExt and MicrOMEGAs (in %), defined as $|\Omega_{\text{RelExt}}-\Omega_{\text{Micro}})|/\Omega_{\text{Micro}}$, as a function of the DM mass. The displayed points have been obtained in a scan with ScannerS which checks for the most relevant theoretical and experimental constraints, for details cf. Refs. [43, 44]. As can be inferred from the plot, for the bulk of the points the difference is below 2%.
@@
Larger differences can occur in models where there is an additional Higgs boson in the visible sector that can have a large total decay width, and for parameter points where we are near an $s$-channel resonance in the (co-)annihilation channel. This is shown in Fig. 4 for the dark doublet phase of the N2HDM [41]. The model, which is based on a 2HDM extended by a real singlet field, features a discrete symmetry such that the Higgs spectrum consists of four dark sector particles (two neutral and two charged ones) and two visible Higgs bosons, one of which is the SM-like Higgs boson. The other non-SM-like Higgs boson, denoted by $h_{\text{non-SM}}$ in the following, can have a large total width, depending on its mass value. The DM particle is given by the lightest dark sector scalar which must be neutral. The figure shows the absolute value of the relative difference between the relic densities calculated with RelExt and MicrOMEGAs, respectively, but now with respect to $m_{h_{\text{non-SM}}}/m_{\text{DM}}$ (left) and $\Gamma_{h_{\text{non-SM}}}$ (right).
@@
Table 3: Parameter point shown in Fig. 5 for the dark doublet phase of the N2HDM. Lines 1-12 contain the input parameters, lines 13-14 the relic densities from RelExt and MicrOMEGAs, respectively. For details about the model and its input parameters, see [41, 42]. For the input parameters with units, all units are in GeV.
@@ A.1.1 The Model CxSM
where $m_{A}$ denotes the DM mass. For the relation between these input parameters and the parameters of the potential, we refer the reader to [39].
@@ A.2.1 The Model TRSM
$V_{\text{Scalar}}=$ $\enspace\mu_{h}^{2}|\Phi_{1}|^{2}+\lambda_{h}|\Phi_{1}|^{4}+m_{1}^{2}\,\phi_{2}^{2}+\frac{\lambda_{2}}{4!}\,\phi_{2}^{4}+m_{2}^{2}\,\phi_{3}^{2}+\frac{\lambda_{3}}{4!}\,\phi_{3}^{4}$
$+$ $\enspace\frac{\lambda_{12}}{2}\,|\Phi_{1}|^{2}\phi_{2}^{2}+\frac{\lambda_{13}}{2}\,|\Phi_{1}|^{2}\phi_{3}^{2}+\frac{\lambda_{\text{23}}}{4}\,\phi_{2}^{2}\phi_{3}^{2}$ (A.17)
$+$ $\enspace m_{12}^{2}\phi_{2}\phi_{3}+\frac{\lambda_{\text{223}}}{4}\,\phi_{2}^{3}\phi_{3}+\frac{\lambda_{\text{233}}}{4}\,\phi_{2}\phi_{3}^{3}+\frac{\lambda_{123}}{2}\,|\Phi_{1}|^{2}\phi_{2}\phi_{3}\,,$
@@ 4 Main M(argv, MODE, BEPS, XTODAY, FAST, CALCWIDTHS, SAVECONTRIBS);
from the \$RelExt/build directory. After increasing the number of scanned points and choosing smaller variations of the couplings for each iteration, the Fig. 4 given in [63] is reproduced.
@@ A.3.1 The Model CPVDM
The model CP in the Dark (CPVDM) [43] is a scalar extension of the SM including two complex doublets $\Phi_{1}$ and $\Phi_{2}$ and one real singlet field $\Phi_{\text{s}}$. In general all the fields can acquire a non-zero vacuum expectation value after spontaneous symmetry breaking. However we will look at the scenario where only $\Phi_{1}$ acquires a neutral non-zero VEV $v$, such that
@@
$V_{\text{Scalar}}=$ $\enspace m_{11}^{2}\Phi_{1}^{\dagger}\Phi_{1}+m_{22}^{2}\Phi_{2}^{\dagger}\Phi_{2}+\dfrac{\lambda_{1}}{2}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\dfrac{\lambda_{2}}{2}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}$
[…]
@@
For a more detailed description on the relations between the potential parameters and the chosen set of input parameters, we refer to [43, 44].
@@ 2 Main M(argv, MODE, BEPS, XTODAY, FAST, CALCWIDTHS, SAVECONTRIBS);
Figure 10: Computed relic density over the DM mass for the CPVDM model. The blue points are generated from a random scan of the best cells on which a random walk was applied. The dark gray points are random points from the entire scan region, while the light gray points are random points from the best cells. The gray dashed line shows the observed relic density. The blue points are compatible with the measured relic density within ±$\pm 2\sigma$, and the red points are validated against the relevant theoretical and experimental constraints using ScannerS.
@@
There is, however, still a benefit in using the random walk below the Higgs threshold. The points obtained during a random walk are not bound to the best cell they originate from. This means that they can produce new best cells from the ones already found. We validated the points obtained from this procedure against the relevant theoretical and experimental constraints using ScannerS [30, 29]. The remaining 214 points are highlighted in red in Fig. 10.
@@
[…]
Random $m_{\text{DM}}<m_{h}$ $10^{5}$ $0.027\%$ $\sim$0.0012 s-1
Best cells $m_{\text{DM}}>m_{h}$ $10^{5}$ $0.29\%$ $\sim$0.0178 s-1
@@ Appendix B Possible Issues with the Generation of FeynRules Model Files
In this section, we describe some possible issues that the user may encounter when implementing a new model in FeynRules. If the model files for FeynRules are already generated, several compatibility checks should be made to ensure that RelExt correctly extracts the relevant information for the relic density calculation. In the following discussion, some of the terminology requires familiarity with the FeynRules language. A detailed description of the program can be found in [28]. The following checks/changes are the most important ones to ensure compatibility with RelExt:
  • In RelExt, the amplitudes relevant for freeze-out will be generated in the Feynman gauge. Some programs such as MadGraph5_aMC@NLO [25] prefer to use the unitary gauge, in which case the Boolean variable \$FeynmanGauge, which usually appears at the top of the model files, will be set to False. In RelExt, \$FeynmanGauge must be set to True to use the Feynman gauge.
@@
[…]
@@ Appendix C Mathematica Code Overview
  • neutraldsmasses is the list with the masses of all neutral dark sector particles.

### v2

@@ 5.4 Boltzmann Equation
After the computation of the TAC, we can move on to the solution of Eq. (5.1). For the effective degrees of freedom $g_{\text{eff}}$, $h_{\text{eff}}$ and $g_{*}^{1/2}$ we use interpolated values from [GONDOLO1991145]. We determine the freeze-out point $x_{\text{f}}$ via the condition given in [GONDOLO1991145],
@@
where $\delta$ can be freely chosen. To find the point $x_{\text{f}}$ which fulfills Eq. (5.13), we use a Bisection method [^9] We only search between $x=5$ and $x=50$. If a point is not found within these boundaries it is not a valid freeze-out point and we return 0 for the relic density. to find the root of the equation. If the option Fast is set to true we use $\delta=1.5$ as recommended in [GONDOLO1991145] and compute the yield via the freeze-out approximation, i.e.
@@
where the $Y_{\text{eq}}^{2}$ term in Eq. (5.1) was discarded. As stated in [Belanger:2004yn] and [GONDOLO1991145], this approximation is accurate within a few percent in almost all scenarios, which we can confirm. When the Fast option is turned off we use $\delta=0.1$ to find the freeze-out point. With this we solve Eq. (5.1) numerically with the initial condition $Y_{\text{f}}=1.1Y_{\text{eq}}(x_{\text{f}})$ at the freeze-out point using an adaptive Dormand-Prince 853 method as described in [10.5555/1403886].
@@ 6 Validation
In Fig. 2, we present the relic density $\Omega_{c}h^{2}$ as a function of the DM mass $m_{\text{DM}}$ for the real singlet model [McDonald:1993ex]. The Higgs sector of this model features a DM particle and a SM-like Higgs with mass 125 GeV. For the plot, we chose the mass range of the DM particle between 5 GeV and 1000 GeV, and the freeze-out portal coupling was set to 0.1 (orange and green points) and 6 (red and black points). We observe, as expected, a resonant annihilation at $m_{\text{DM}}=m_{h}/2$, where $m_{h}$ denotes the SM-like Higgs mass value of 125 GeV, resulting in a dip in the relic density. The subsequent peaks and dips in the relic density are a consequence of the DM mass passing the thresholds for on-shell production of, respectively, a $W$ boson, $Z$ boson and finally a SM-like Higgs pair. The dip at the Higgs-mass threshold is much less pronounced for $\lambda_{\text{portal}}=0.1$ than for $\lambda_{\text{portal}}=6$, reflecting the different values of the SM-like Higgs coupling to the DM particles in the two chosen scenarios. As can be inferred from the figure, the MicrOMEGAs and the RelExt results are in very good agreement, with differences of at most 6%.
@@
Next we compare results in the model CP in the Dark [Azevedo:2018fmj, Biermann:2022meg]. It is based on an N2HDM with an imposed discrete symmetry such that the Higgs spectrum consists of a SM Higgs boson and the additional Higgs bosons, three neutral and two charged ones, residing in a dark sector with explicit CP violation. The lightest of these dark scalars must be neutral and represents the DM candidate. In Fig. 3, we show the absolute value of the relative difference in the relic density between RelExt and MicrOMEGAs (in %), defined as $|(\Omega_{\text{RelExt}}-\Omega_{\text{Micro}})|/\Omega_{\text{Micro}}$, as a function of the DM mass. The displayed points have been obtained in a scan with ScannerS which checks for the most relevant theoretical and experimental constraints, for details cf. Refs. [Azevedo:2018fmj, Biermann:2022meg]. As can be inferred from the plot, for the bulk of the points the difference is below 2%.
@@
Larger differences can occur in models where there is an additional Higgs boson in the visible sector that can have a large total decay width, and for parameter points where we are near an $s$-channel resonance in the (co-)annihilation channel. This is shown in Fig. 4 for the dark doublet phase of the N2HDM [Engeln:2020fld]. The model, which is based on a 2HDM extended by a real singlet field, features a discrete symmetry such that the Higgs spectrum consists of four dark sector particles (two neutral and two charged ones) and two visible Higgs bosons, one of which is the SM-like Higgs boson. The other non-SM-like Higgs boson, denoted by $h_{\text{non-SM}}$ in the following, can have a large total width, depending on its mass value. The DM particle is given by the lightest dark sector scalar which must be neutral. The figure shows the absolute value of the relative difference between the relic densities calculated with RelExt and MicrOMEGAs, respectively, but now with respect to $m_{h_{\text{non-SM}}}/m_{\text{DM}}$ (left) and $\Gamma_{h_{\text{non-SM}}}$ (right).
@@
Table 3: Parameter point shown in Fig. 5 for the dark doublet phase of the N2HDM. Lines 1-12 contain the input parameters, lines 13-14 the relic densities from RelExt and MicrOMEGAs, respectively. For details about the model and its input parameters, see [Engeln:2020fld, Azevedo:2021ylf]. For the input parameters with units, all units are in GeV.
@@ A.1.1 The Model CxSM
where $m_{A}$ denotes the DM mass. For the relation between these input parameters and the potential parameters,reader to [Egle:2022wmq].
@@ A.2.1 The Model TRSM
$V_{\text{Scalar}}=$ $\kern 5.0pt\mu_{h}^{2}|\Phi_{1}|^{2}+\lambda_{h}|\Phi_{1}|^{4}+m_{1}^{2}\,\phi_{2}^{2}+\frac{\lambda_{2}}{4!}\,\phi_{2}^{4}+m_{2}^{2}\,\phi_{3}^{2}+\frac{\lambda_{3}}{4!}\,\phi_{3}^{4}$
$+$ $\kern 5.0pt\frac{\lambda_{12}}{2}\,|\Phi_{1}|^{2}\phi_{2}^{2}+\frac{\lambda_{13}}{2}\,|\Phi_{1}|^{2}\phi_{3}^{2}+\frac{\lambda_{\text{23}}}{4}\,\phi_{2}^{2}\phi_{3}^{2}$ (A.17)
$+$ $\kern 5.0ptm_{12}^{2}\phi_{2}\phi_{3}+\frac{\lambda_{\text{223}}}{4}\,\phi_{2}^{3}\phi_{3}+\frac{\lambda_{\text{233}}}{4}\,\phi_{2}\phi_{3}^{3}+\frac{\lambda_{123}}{2}\,|\Phi_{1}|^{2}\phi_{2}\phi_{3}\,$,
@@ 4 Main M(argv, MODE, BEPS, XTODAY, FAST, CALCWIDTHS, SAVECONTRIBS);
from the \$RelExt/build directory. After increasing the number of scanned points and choosing smaller variations of the couplings for each iteration, the Fig. 4 given in [Capucha:2024oaa] is reproduced.
@@ A.3.1 The Model CPVDM
The model CP in the Dark (CPVDM) [Azevedo:2018fmj] is a scalar extension of the SM including two complex doublets $\Phi_{1}$ and $\Phi_{2}$ and one real singlet field $\Phi_{\text{s}}$. In general all the fields can acquire a non-zero vacuum expectation value after spontaneous symmetry breaking. However we will look at the scenario where only $\Phi_{1}$ acquires a neutral non-zero VEV $v$, such that
@@
$V_{\text{Scalar}}=$ $\kern 5.0ptm_{11}^{2}\Phi_{1}^{\dagger}\Phi_{1}+m_{22}^{2}\Phi_{2}^{\dagger}\Phi_{2}+\dfrac{\lambda_{1}}{2}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\dfrac{\lambda_{2}}{2}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}$
$+$ $\kern 5.0pt\lambda_{3}\Phi_{1}^{\dagger}\Phi_{1}\Phi_{2}^{\dagger}\Phi_{2}+\lambda_{4}\Phi_{1}^{\dagger}\Phi_{2}\Phi_{2}^{\dagger}\Phi_{1}+\dfrac{\lambda_{5}}{2}\left[\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+\text{h.c.}\right]$ (A.23)
$+$ $\kern 5.0pt\dfrac{1}{2}m_{s}^{2}\Phi_{\text{s}}^{2}+\dfrac{\lambda_{6}}{8}\Phi_{\text{s}}^{4}+\dfrac{\lambda_{7}}{2}\Phi_{1}^{\dagger}\Phi_{1}\Phi_{\text{s}}^{2}+\dfrac{\lambda_{8}}{2}\Phi_{2}^{\dagger}\Phi_{2}\Phi_{\text{s}}^{2}+(A\Phi_{1}^{\dagger}\Phi_{2}\Phi_{s}+\text{h.c.})\,$,
@@
For a more detailed description on the relations between the potential parameters and the chosen set of input parameters, we refer to [Azevedo:2018fmj, Biermann:2022meg].
@@ 2 Main M(argv, MODE, BEPS, XTODAY, FAST, CALCWIDTHS, SAVECONTRIBS);
Figure 10: Computed relic density over the DM mass for the CPVDM model. The blue points are generated from a random scan of the best cells on which a random walk was applied. The dark gray points are random points from the entire scan region, while the light gray points are random points from the best cells. The gray dashed line shows the observed relic density. The blue points are compatible with the measured relic density within $\pm 2\sigma$, and the red points are validated against the relevant theoretical and experimental constraints using ScannerS.
@@
There is, however, still a benefit in using the random walk below the Higgs threshold. The points obtained during a random walk are not bound to the best cell they originate from. This means that they can produce new best cells from the ones already found. We validated the points obtained from this procedure against the relevant theoretical and experimental constraints using ScannerS [Muhlleitner:2020wwk, Coimbra:2013qq]. The remaining 214 points are highlighted in red in Fig. 10.
@@
Random $m_{\text{DM}}>m_{h}$ $10^{5}$ $0.\%$ $\sim$0 s-1
Random $m_{\text{DM}}<m_{h}$ $10^{5}$ $027\% \sim0.0$012 s-1
Best cells $m_{\text{DM}}>m_{h}$ $10^{5}$ $29\% \sim0.0$178 s-1
@@ Appendix B Possible Issues with the Generation of FeynRules Model Files
In this section, we describe some possible issues that the user may encounter when implementing a new model in FeynRules. If the model files for FeynRules are already generated, several compatibility checks should be made to ensure that RelExt correctly extracts the relevant information for the relic density calculation. In the following discussion, some of the terminology requires familiarity with the FeynRules language. A detailed description of the program can be found in [Alloul:2013bka]. The following checks/changes are the most important ones to ensure compatibility with RelExt:
  • In RelExt, the amplitudes relevant for freeze-out will be generated in the Feynman gauge. Some programs such as MadGraph5_aMC@NLO [Alwall:2011uj] prefer to use the unitary gauge, in which case the Boolean variable \$FeynmanGauge, which usually appears at the top of the model files, will be set to False. In RelExt, \$FeynmanGauge must be set to True to use the Feynman gauge.
@@
  • Particle classes can be used to specify the symbol and the numerical value for the masses of the different members of a particle class using the Mass option, as in Mass $\to$ {MH, Internal}. In this example, MH is defined as an internal parameter (even though it is a free parameter), and the user needs to define MH again in the M\$Parameters list.
@@ Appendix C Mathematica Code Overview
  • neutraldsmasses is the list with the masses of all neutral dark sector particles. d